OpenStudy (anonymous):

(x+4)^2+(y-3)^2=25 domain= range= function?= x-intercepts= y-intercepts= center= radius=

OpenStudy (anonymous):

looks like a circle right? do you know the center and the radius?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

\[(x-h))^2+(y-k)^2=r^2\] is a circle with center \((h,k)\) and radius \(r\)

OpenStudy (anonymous):

in your case \(h=-4,k=3\) and \(r^2=25\)

OpenStudy (anonymous):

here is a nice picture http://www.wolframalpha.com/input/?i=circle+%28x%2B4%29^2%2B%28y-3%29^2%3D25

OpenStudy (anonymous):

what are the x and y intercepts?

OpenStudy (anonymous):

@satellite73